OpenStudy (anonymous):

Solve (3y+2x-8)dx = (x+2y-3)dy

OpenStudy (anonymous):

u want me to factor? or solve for d and y?

OpenStudy (anonymous):

i think this is ur answer: −2dxy+2dy2−3dy=2dx2−8dx

OpenStudy (anonymous):

this is how u get it

OpenStudy (anonymous):

(3y+2x−8)dx=(x+2y−3)dy Since y is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation. dy(x+2y−3)=dx(3y+2x−8) Multiply dx by each term inside the parentheses. dy(x+2y−3)=3dxy+2dx2−8dx Multiply dy by each term inside the parentheses. dxy+2dy2−3dy=3dxy+2dx2−8dx Since 3dxy contains the variable to solve for, move it to the left-hand side of the equation by subtracting 3dxy from both sides. dxy+2dy2−3dy−3dxy=2dx2−8dx Since dxy and −3dxy are like terms, add −3dxy to dxy to get −2dxy. −2dxy+2dy2−3dy=2dx2−8dx

OpenStudy (anonymous):

@dorin2